# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v2

@@ 3 Methods
This study specifically investigates recurrent neural networks for virtual analog modeling. In particular, we aim to identify which type of recurrent layer or architectural configuration provides accurate modeling of different types of audio effects, if any. These architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states, rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks, which present low computational complexity and low input-output, allowing implementation on real-time consumer-grade digital audio systems. We have developed models that exhibit minimal input-output delay and are well-suited for the typical audio processing workflow that operates on data in blocks. The designed architectures primarily consist of a recurrent layer along with a conditioning block. This architectural choice helps us incorporate control parameters. The designed models are applied to different audio effects and finally evaluated using different metrics. The primary objective of this study is to explore the capabilities of the different recurrent models in learning from raw digital audio and effectively modeling audio effects. These insights can be used to understand the goodness and weakness of these models and which can suit better each audio effect. The models here presented, having limited parameters, are not expected to lead to perfectly accurate emulation, but a good starting point.
After selecting several types of analog audio effects to be included in this comparative study, as previously discussed, we have looked for existing public datasets including variable control parameters. In case these are not available, we recorded the dataset using analog effects within our availability. For those not available, we built the dataset using software plugins emulating specific analog effects. Datasets, for hardware or software effects, were collected using a tool that we have developed for this specific purpose 111https://github.com/stefanofasciani/DGMD.
In the rest of this section, we first detail the four architectures used in this comparative study, then we present the the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), L-ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. L-ED is based on the concept of internal state sharing [11]. All of these architectures follow the same design, consisting of recurrent layer, layer normalization, conditioning layer, and skip connections.
The number of input samples is identical for all architectures and set to . This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below ms (i.e., samples at kH.). The input of all architectures is an array including the most recent input sample, used to generate the single current output sample. As a consequence, the model is required to perform predictions at audio rate, one for each output audio sample. The models map a vector of sample in one scalar output sample:
where is the discrete time index and is the overall function describing the models. Before being fed into the recurrent layer (LSTM, LRU, or S4D), the input samples array is processed by a fully connected layer that performs a linear projection:
@@
The first two architectures use LSTMs as the recurrent layer, which was designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information. The LSTM architecture is governed by the following equations:
@@
In the case of LSTM-ED, the input vector is equally split into two vectors, containing and containing containing . The first vector is fed to an LSTM layer, which computes the internal states, , that serve as initial states for a second LSTM layer representing the actual recurrent layer. The second part of the input vector is linearly transformed and fed into the second LSTM, following the same procedure as the other models.
The third architecture is based on LRU. Compared to a vanilla RNN, the LRU uses a linear activation function used to compute the hidden state, instead of a nonlinear function. This modification allows the memory capability to improve significantly. The LRU equations are the following:
where is the hidden state at time and acts as ’memory’ of the network, the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization, and an exponential parameterization, while a normalization scheme for the hidden states is used to boost the efficiency and accuracy [7].
@@
The remarkable ability of SSMs to capture long-range dependencies stems from their utilization of a specific state matrix known as the ”HiPPO matrix” [58]. This matrix is designed to encode all the past input history in , finding a map from to a higher dimensional space that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state encodes the history of the signal. In our study, we employ the diagonal state space model variant, called S4D model [59], which is defined by parameterizing its state matrix as a diagonal matrix, boosting the efficiency of the model. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials , resulting in the SSM decomposing the input signal onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure. This characteristic gives the SSM the ability to effectively model long-range dependencies.
The output size of the recurrent layer will be equal to the number of hidden units of the layer. In the case of LRU and S4D a fully connected layer with a softsign activation function also follows the recurrent layer, to compensate for a lack of nonlinear activation functions in these layers:
Subsequently, the Layer Normalization [60] is applied to help the training process. Layer Normalization estimates mean and variance which is then used to normalize the summed input to the weights, of the previously hidden layer on each training case. The mean and the standard deviation are computed over all the weights of the previous layer to re-center and re-scale :
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Initial experiments indicate improved performance with the conditioning block is placed after the recurrent layers rather than before. This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output, rather than influencing the inference of the recurrent layer based on the control parameters.
Finally, a skip connection is applied to the output of the output layer, which is a fully connected layer with one unit, , where the input sample corresponding to the same time index is added to :
@@
The number of units is chosen to have approximately trainable parameters among the architectures and effects since an identical number is achievable due to major differences in the layers. As a result, the selected LSTM and LSTM-ED present units in the recurrent layer and in fully connected ones. LRU and S4D present units in the recurrent layer, and in fully connected ones. The number of trainable parameters also depends on the dimensional of the conditioning input , equal to the number of variable parameters in the datasets.
@@ 3.2 Losses and Metrics
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, and represent the target and predicted signal. The metrics compare and using non-overlapping segments of the output signal with size samples. The symbol represents the overall number of samples in the output signal.
@@
The losses and metrics presented so far, calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric that considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library 222https://github.com/magenta/ddsp developed after [1]:
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to and hop size equal to samples. In the formula, represents the hop size. This metric emphasizes the ability to accurately predict portions of the output sound with rapid changes.
@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. As already mentioned, in our search, we sought out publicly available datasets encompassing all the aforementioned types of analog audio effects and considering the presence of adjustable control parameters, if any. If we couldn’t find these datasets online, we proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 1 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box [49], [48]. For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal 333https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron 444https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator 555https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech 666http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA2A 777https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins: the Helper Saturator 888https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ 999https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
The CL 1B and the LA2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The amplitude of the input signal determines the brightness of the element that, in turn, changes the resistance in the gain attenuation circuit. We used the available datasets [50] [62]. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. To limit the data, we consider the threshold values range from to dBu, ratio values from : to :, attack time from to ms and release time from to seconds. The total number of combinations was . In addition, we used the of each recording. On the other hand, LA2A presents the peak reduction knob indicating the amount of compression ([0,100]) and a mode switch (limiter, compression), resulting in combinations. Similarly to the CL1B dataset, we consider of each recording.
@@
LA2A111Hardware. 333Downloaded. | Optical | Peak Reduction | [0,100] | |
@@ 3.3.1 Data Collection
The input mono signal differs if the device being measured includes control parameters or not. In particular, if any, the signal has a duration of seconds for each parameter combination. If do not present any control parameter, to the signal is added other sound portions to achieve a duration of minutes. The input signals include a variety of sounds, such as frequency sweeps covering a range of Hz to kHz, white noises with increasing amplitudes (both linear and logarithmic), recordings of instruments such as guitar, bass, drums (both loops and single notes), vocals, piano, pad sounds, and sections from various electronic and rock songs. Finally, the control parameter of each effect was mapped to the range of [0,1].
@@ 3.4 Experimenting and Learning
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
The composition of the training, validation, and test sets varies across datasets. For publicly available datasets such as LA2A, the three sets are already split and we utilized the sets in their original form. However, for the dataset we recorded, we used for the training set, for the validation set, and the remaining for the test set. For effects with variable control parameters, each s recording in the dataset is associated with a specific combination of the effect parameters. The split was carried out at the individual recording level, which ensures that all combinations of control parameters present an equal share in each subset, and are equally used for training and evaluation of the model. The initial of all the recordings are used as the test set. The following portion of the recordings is used as a validation set, while the remaining parts were used as the training set. Minor manual adjustments are made to ensure that splitting points fall within segments of silence.
@@ 4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 3, and 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 2. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging, based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure 3, the spectral flux in Figure 4, the spectra in Figure 5, and the spectrograms in Figure 6. For datasets including conditioning parameters, the plots showing predictions versus the target, are generated for a seconds input signal containing bass, guitar, and drum loop sound. For the dataset without conditioning, the plots are generated from a seconds input signal with percussive sounds. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first seconds.
@@
In the OD300 overdrive dataset, the LSTM model presents the lowest MSE, MAE, and ESR, indicating better general amplitude matching between target and prediction. The S4D model is also accurate, especially when considering the transients and the RMS envelope. This can be observed from the lower error in terms of RMSE and spectral flux. On the other hand, in the case of maximum distortion, as seen in Figures 3 and 4, this superiority is less perceived. LSTM-ED and LRU are left behind except when looking where LSTM-ED show better performance than LSTM. Notably, the spectrograms shown in Figure 6 illustrate that S4D model generates more harmonics in the higher frequency region than LSTM, and especially than LSTM-ED and LRU whose produce a cleaner signal. This could be perceptually advantageous compared to the other model that poorly predicts such region of the spectrum, as summarized by the errors. From Figure 5 it is evident how the LSTM-based produces the best matching for energy at low frequencies. All models except LRU hit the early stop training condition at a relatively close epoch.
@@
The Neutron OD dataset shows larger errors and it appears to be a more challenging modeling case for these architectures. Both LSTM and S4D models yield similar results in terms of average amplitude and frequency errors. S4D show slightly better performance. LSTM-ED surpasses the previous models in terms of RMSE, NESR, and SF, despite exhibiting a larger RMSE mismatch at maximum distortion settings, as visible in Figure 3. The spectrograms show less difference in harmonic distortion generation, as reflected in the frequency errors, which are very close to each other, although the S4D model exhibits lower errors. When informally listening, S4D tends to add more distortion than needed, while LSTM-based models do the opposite. Also in this case, LRU presents an early stopping epoch quite different than the other three cases.
For both the BBD and Saturator datasets, S4D outperforms all other models, while LSTM-ED is slightly less accurate. This is reflected also in the Figures 3 and 4. The LSTM and LRU models exhibit artifacts, as evident in Figure 6, which negatively impact performance, especially for LRU, and can be clearly heard when listening to the predicted output audio examples. The BBD dataset does not consider conditioning, unlike the saturator dataset, making it a simpler task, producing no significant deviation among the models’ performances. The LSTM model tends to overpredict the energy at lower frequencies and underpredicts at higher frequencies, whereas the LRU model significantly underpredicts the signal amplitude compared to the other models. When informally listening, S4D model is the only one controlling the amount of saturation. The other models tend to overcharge the saturation.
In the case of the equalizer, LSTM-based models, specifically LSTM and LSTM-ED, exhibit slightly better performance compared to LRU and S4D. This holds also in the challenging conditioning scenario as visible in Figures 3, 4, 6, and 5.
The low-pass filter is relatively more difficult than the equalizer, as it introduces more drastic changes in the frequency contents of the processed signal. As the early stopping condition is met after a very few epochs, it appears that the model training converges quickly, as visible in Table 2. However, when informally listening to the predicted output, it is evident that all models fail in learning the filter behavior as they simply attenuate the signal rather than removing high-frequency contents. Furthermore, the LSTM model produces better results when considering the time-domain metrics, but the errors are similar to other models when considering the frequency-domain ones. Also, the model tends to predict an excessive amount of energy in the lower portion of the spectrum, as visible in Figure 5. When considering the extreme conditioning case S4D performs better when considering RMSE, the transients, and spectrogram, as visible Figures 3 4 and 6.
@@
Helper | LSTM | BBD | LSTM | ||||
Saturator | L-ED | Analog | L-ED | ||||
LRU | Delay | LRU | |||||
S4D | S4D | ||||||
Pultec | LSTM | Neutron’s | LSTM | ||||
Passive | L-ED | Filter | L-ED | ||||
EQ | LRU | LRU | |||||
S4D | S4D | ||||||
LA2A | LSTM | CL1B | LSTM | ||||
L-ED | L-ED | ||||||
@@
The optical compressor exhibits a time-variant response, and especially prolonged temporal dependencies dictated by the variable parameters of attack and release. This is particularly evident in the CL1B compressor, which poses a more challenging modeling task for neural networks. The LSTM model is the less accurate in this case. It presents a larger loss in both compressor cases. For the LA2A case, we can see a significant mismatch for RMSE and in transients phases, in Figure 3 and 4. In terms of the frequency domain, S4Dprovides a better match overall, although the LSTM model performs better in terms of MFCC error. Indeed spectrograms associated with LSTM-based models appear more similar to the target output than those associated with the S4D model. In the case of the compressor, the ability to encode past information is crucial, and S4D exhibits superior accuracy. The L-ED model is designed to improve the temporal tracking abilities of basic LSTM. This fact is confirmed by the results for both compressor devices. This improvement is particularly evident with the CL1B when setting long attack and release time through the conditioning control parameters. These reach up to ms and seconds, respectively. In Figure 3 and 4 difference across models is less marked, but the superiority of S4D is still clear. LRU presents smaller errors than LSTM as well but it produces clearly audible artifacts, as visible in Figure 6.
Generally, for most effects considered in this study, performance metrics are fairly close across models. Although S4D outperforms other models by a small extent, considering the metrics only, it seems to be the most suitable architecture, among those included in this study, for generalizing the problem of analog effect modeling. LSTM-ED and in particular S4D are more efficiently encoding, inside the states, past signal information. Indeed in the case of optical compressors, they outperform other models significantly. On the other hand, LSTM appears to perform well only when emulating effects where time-dependencies are shorter and less crucial, such as in the case of overdrive, equalizer, and filter. In addition, LSTMs tend to predict more accurately the energy on the low-frequency components. Finally, LRU seems to be not appropriate for modeling audio effects as they introduce major audible artifacts.

### v3

@@ 3 Methods
This study specifically investigates recurrent neural networks for virtual analog modeling. In particular, we aim to identify which type of recurrent layer or architectural configuration provides accurate modeling of different types of audio effects, if any. These architectures offer advantages due to their ability to capture time dependencies in the data by utilizing internal states rather than solely relying on the input received at each iteration. In addition, we restrict the experimentation to relatively small networks, which present low computational complexity and low input-output, allowing implementation on real-time consumer-grade digital audio systems. We have developed models that exhibit minimal input-output delay and are well-suited for the typical audio processing workflow that operates on data in blocks. The designed architectures primarily consist of a recurrent layer along with a conditioning block. This architectural choice helps us incorporate control parameters. The designed models are applied to different audio effects and finally evaluated using different metrics. The primary objective of this study is to explore the capabilities of the different recurrent models in learning from raw digital audio and effectively modeling audio effects. These insights can be used to understand the goodness and weakness of these models and determine which can better suit each audio effect. The models presented here, having limited parameters, are not expected to lead to perfectly accurate emulation but a good starting point.
After selecting several types of analog audio effects to be included in this comparative study, we recorded the dataset using analog effects within our availability. We built the dataset using software plugins emulating specific analog effects. Datasets for hardware or software effects were collected using a tool that we have developed for this specific purpose 111https://github.com/stefanofasciani/DGMD.
In the rest of this section, we first detail the four architectures used in this comparative study; then we present the the metrics we used for evaluating the trained models and the specific audio effects from which we collected the datasets. Finally, explain the process we used to train and test the models.
@@ 3.1 Architectures
The following architectures, based on different types of recurrent layers, are compared in this study: LSTM (a), ED (b), LRU (c), and S4D (d). The four architectures are illustrated in Figure 1. ED is based on the concept of internal state sharing [11]. These architectures follow the same design, consisting of a linear, fully connected layer, a recurrent layer, and a conditioning block.
The number of input samples is identical for all architectures and set to . This choice represents a trade-off: on the one hand, there is the amount of past input information used to generate the current output, where more is typically better; on the other hand, there is the resulting input-output latency, which increases with the number of input samples. Our goal is to keep this latency below ms (i.e., samples at kH.). The input of all architectures is an array, including the most recent input sample, used to generate the single current output sample. Consequently, the model must perform predictions at an audio rate, one for each output audio sample. The models map a vector of sample in one scalar output sample:
where is the discrete time index, and is the overall function describing the models. Before being fed into the recurrent layer (LSTM, LRU, or S4D), the input samples array is processed by a fully connected layer that performs a linear projection:
@@
The first two architectures use LSTMs as the recurrent layer, designed to improve the memory capacity of the vanilla RNN by incorporating internal mechanisms known as gates, regulating the flow of information. The LSTM architecture is governed by the following equations:
@@
In the case of ED, the input split into two equal-size parts, containing and containing containing . The first part feds two separate convolutional layers, which compute the states, . These states are combined with the internal states of the LSTM layer computed during the previous iteration. A sigmoid function is applied to the LSTM internal states and then element-wise multiplied to . This operation determines how much information should be taken from the computed . The result of this operation represents the new LSTM’s internal states. This process is detailed in Figure 2.
The second part of the input vector is linearly transformed and fed into the second LSTM, following the same procedure as the other models.
The third architecture is based on LRU. Compared to a vanilla RNN, the LRU uses a linear activation function to compute the hidden state instead of a nonlinear function. This modification allows the memory capability to improve significantly. The LRU equations are the following:
where is the hidden state at time and acts as ’memory’ of the network, the output vector. LRU keeps the recurrent layer linear but moves the nonlinearity to a subsequent fully connected layer. Additionally, to further enhance stability, the LRU incorporates a complex-valued diagonal recurrent matrix initialization and an exponential parameterization. At the same time, a normalization scheme for the hidden states is used to boost the efficiency and accuracy [7].
@@
The remarkable ability of SSMs to capture long-range dependencies stems from their utilization of a specific state matrix known as the ”HiPPO matrix” [58]. This matrix is designed to encode all the past input history in , finding a map from to a higher dimensional space that represents the compression of the history. The matrix allows the model to be conceptualized as a convolutional model that decomposes an input signal onto an orthogonal system of smooth basis functions. In this way, the state encodes the history of the signal. In our study, we employ the diagonal state space model variant, S4D model [59], which is defined by parameterizing its state matrix as a diagonal matrix, boosting the model’s efficiency. By doing so, the basis kernels have closed-form formulas represented by normalized Legendre polynomials , resulting in the SSM decomposing the input signal onto an infinitely long set of basis functions that are orthogonal with respect to an exponentially decaying measure. This characteristic gives the SSM the ability to model long-range dependencies effectively.
The recurrent layer’s output size will equal the number of hidden units of the layer. In the case of LRU and S4D, a fully connected layer with a hyperbolic tangent activation function also follows the recurrent layer to compensate for a lack of nonlinear activation functions in these layers:
The conditioning block is then applied just before the output layer, which consists of a linear fully connected layer with one unit, which is described in the following subsection. Initial experiments indicate improved performance when the conditioning block is placed after the recurrent layers rather than before. This suggests that it is more beneficial for the networks to use the information given by the control parameters to project the output of the recurrent layer and determine the extent to which this information influences the final output rather than influencing the inference of the recurrent layer based on the control parameters.
Finally, the output layer, which is a fully connected layer with one unit, computes :
@@
The number of units is chosen to have approximately trainable parameters among the architectures and effects since an identical number is achievable due to major differences in the layers. As a result, the selected LSTM and ED present units in the recurrent layer and in fully connected ones. LRU and S4D present units in the recurrent layer, and in fully connected ones. The number of trainable parameters also depends on the dimensional of the conditioning input , equal to the number of variable parameters in the datasets.
@@ 3.2 Losses and Metrics
In addition, our quantitative evaluation is based on various metrics to assess the models from different perspectives. In the following expressions, and represent the target and predicted signal, and the symbol represents the overall number of samples in the output signal.
@@
The losses and metrics presented so far calculate the average difference between the target and prediction, which may hide large errors that occur only during signal transients. Transients, such as sharp attacks or decays, may represent a small fraction of the overall duration of an audio signal. In this case, we may obtain a relatively low error even with all transients poorly modeled. Therefore, we have included a metric that considers the differences between pairs of spectrums computed on consecutive overlapping windows, as used in the library 222https://github.com/magenta/ddsp developed after [1]:
where the Short-Time Fourier Transform (STFT) is computed with a window size equal to and hop size equal to samples. This metric emphasizes the ability to accurately predict portions of the output sound with rapid changes.
@@ 3.3 Audio Effects and Datasets
In this section, we introduce the audio effects selected for the study and the composition of their datasets. We proceeded to record the dataset using effects that were at our disposal. In particular, we selected the following effects: overdrive, saturation, equalization, low-pass filter, and compression. Table 1 summarizes all the datasets included in the study. Delay-based effects require a different approach than a direct black box [49], [48]. For this reason, they are not included in this study.
The dataset we use is derived from the following hardware devices: Behringer OD300 overdrive pedal 333https://www.behringer.com/product.html?modelCode=P0608, Behringer Neutron 444https://www.behringer.com/product.html?modelCode=P0CM5’s overdrive module, Behringer Neutron’s filter module in low-pass mode, TC Electronic Bucket Brigade Analog Delay used as a saturator 555https://www.tcelectronic.com/product.html?modelCode=P0EBV, CL 1B TubeTech 666http://www.tube-tech.com/cl-1b-opto-compressor/and Teletronix LA-2A 777https://www.uaudio.com/hardware/la-2a.html optical compressors. We have also collected data from two software VA plugins: the Helper Saturator 888https://www.waproduction.com/plugins/view/helper-saturator and Universal Audio Pultec EQ 999https://www.uaudio.com/uad-plugins/equalizers/pultec-passive-eq-collection.html, because the hardware saturator presents no variable control parameter, and because we have no analog equalizer in our immediate availability.
@@
The CL 1B and the LA-2A are analog optical compressors. In optical compressors, a lighting-emitting element is fed with the audio signal that illuminates a light-sensitive resistor. The input signal’s amplitude determines the element’s brightness, which, in turn, changes the resistance in the gain attenuation circuit. We recorded the dataset directly from the compressor. The CL 1B compressor dataset presents four parameters: attack, release, ratio, and threshold. We consider the threshold values to range from to dBu, ratio values from : to :, attack time from to ms, and release time from to seconds. We consider three equally spaced values for the ratio, attack, and release time and four values for the threshold. The resulting number of parameter combinations is . On the other hand, LA-2A presents the peak reduction knob indicating the amount of compression ([,]) and a mode switch (limiter, compression). We consider ten equally spaced values for the peak reduction, resulting in parameter combinations.
@@
| LA-2A111Hardware. 333Downloaded. | Optical | Peak Reduction | [0,100] | |
@@ 3.3.1 Data Collection
The input mono signal has a duration of seconds for each parameter combination and includes a variety of sounds, such as frequency sweeps covering a range of Hz to kHz, white noises with increasing amplitudes (both linear and logarithmic), recordings of instruments such as guitar, bass, drums (both loops and single notes), vocals, piano, pad sounds, and sections from various electronic and rock songs. Finally, the control parameter of each effect was mapped to the range of [0,1].
@@ 3.4 Experimenting and Learning
We adopted the same learning schedule and minibatch size for all models across all datasets. As explained in Section 3.3, we collected the datasets using the same approach and the same input file to minimize the difference among them. However, it is important to highlight that each dataset and modeling task may benefit from specific adjustments to achieve more accurate results. Different datasets and tasks may require slightly different learning rate schedules; an excessively slow or fast learning rate can lead to suboptimal solutions. The number of samples to process before updating the weights, referred to as the minibatch size, impacts the frequency of updating weights: larger sizes may result in faster convergence but poorer generalization, while smaller sizes can aid generalization but may require additional epochs.
The composition of the training, validation, and test sets is identical across datasets. We used for the training set, for the validation set, and the remaining for the test set. For effects with variable control parameters, each s recording in the dataset is associated with a specific combination of the effect parameters. The split was carried out at the individual recording level, ensuring that all control parameter combinations present an equal share in each subset and are equally used for training and evaluation of the model. The initial of all the recordings are used as the test set. The following portion of the recordings is used as a validation set, while the remaining parts were used as the training set. Minor manual adjustments are made to ensure that splitting points fall within segments of silence.
@@ 4 Results
The evaluation is based on the metrics described in Section 3.2 and presented in Table 3, and 4. The MSE loss and the epoch-related epoch resulting in lower validation loss during training are reported in Table 2. The losses and the metrics consider all conditioning scenarios and provide an average error across all combinations of parameters. To further verify the model’s capacity to learn the conditioned effect, we provide a series of plots that compare the target signal with the predictions generated by the four distinct architectures. For these plots, we chose the modeling scenario that we believe to be the most challenging based on the effect parameters used as conditioning factors for the networks. Except for the low-pass filter, this typically involves parameters at the maximum of the selected range since the output signal, under these conditions, is likely to deviate most significantly from the input. Regarding the low-pass filter, the parameters are configured for minimum cutoff and maximum resonance. The plots show the RMSE in Figure 4, the spectral flux in Figure 5, the spectra in Figure 7, and the spectrograms in Figure 6. For datasets including conditioning parameters, the plots showing predictions versus the target are generated for a seconds input signal containing bass, guitar, and drum loop sound. These input signals are part of the respective test sets. To ensure a clearer comparative visualization of the spectral flux, we use only the first seconds.
@@
In the OD300 overdrive dataset, the LSTM model presents the lowest MSE, MAE, and ESR, indicating better general amplitude matching between target and prediction. The S4D model is also accurate, especially when considering the transients and the RMS envelope. This can be observed from the lower error in terms of RMSE and spectral flux. On the other hand, in the case of maximum distortion, as seen in Figures 4 and 5, this superiority is less perceived. ED and LRU are left behind except when looking where ED show better performance than LSTM. Notably, the spectrograms shown in Figure 6 illustrate that S4D model generates more harmonics in the higher frequency region than LSTM, and especially than LRU whose produces a cleaner signal. This could be perceptually advantageous compared to the other model that poorly predicts such region of the spectrum, as summarized by the errors. From Figure 7 it is evident how the LSTM-based produces the best matching for energy at low frequencies.
@@
Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the architectures being investigated. Both LSTM and LRU models yield similar results in terms of average amplitude and frequency errors. The LSTM models show slightly better performance in the time domain, while the LRU excel in the frequency domain. As observed in Figure 4, both LSTM and LRU model achieves a better match with the target. However, the LSTM model tends to produce more energy, whereas the LRU model yields less energy than the target. The S4D model surpasses previous models in terms ofSF, despite exhibiting a larger mismatch at maximum distortion settings, as visible in Figure 5. The spectrograms indicate that the S4D model exhibits a slightly better match than the LSTM model, even though the latter exhibits smaller errors. Upon informal listening, the S4D model tends to add less distortion necessary, whereas the LSTM-based models tend to do the opposite. Lastly, the LRU model exhibits a later stopping epoch than the other three cases.
For models trained on the Saturator dataset, the LRU model presents lower MSE, MAE, and values. Meanwhile, the S4D model outperforms the others with respect to ESR and metrics. The ED model is slightly less accurate but performs similarly to S4D. In terms of frequency domain errors, LRU also produces better results. As evidenced in Figures 4 and 5, the S4D and ED models achieve closer matching to the target, whereas the LRU model exhibits significantly lower energy, and the LSTM model introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance. These artifacts can also be seen in Figure 5 and 6. Despite these issues, the S4D and ED models still provide a better match with the target spectrograms. The same trend is observed in Figure 7, suggesting that the LRU model performs better in scenarios where minimal sound alteration is required.
When training models with the Equalizer dataset, the LRU and S4D models, which are both LSTM-based, exhibit slightly better performance compared to the LSTM and ED models. In particular, the LSTM model performs poorly in the frequency domain. The S4D model shows improvements in and metrics. This enhanced performance is also evident in challenging conditioning scenarios, as visible in Figures 4, 5, 6, and 7. The LSTM model once again exhibits some audible artifacts, as can be seen in Figure 5.
The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters. This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling. Among the tested models, the LSTM is the least accurate, exhibiting higher loss in both the LA-2A and CL 1B cases. In contrast, the S4D and ED models demonstrate superior accuracy for these tasks. Specifically, S4D outperforms the ED in the LA-2A case, while the textbfED excels in the CL 1B case when evaluated using time domain-based metrics. However, the trend reverses in the frequency domain, although ED consistently shows lower and values for both cases.
In the LA-2A case, we observe an overall good match for RMSE in Figure 4 and for transient phases in Figure 5. The LSTM and LRU models display slightly more mismatches in both plots. Upon examining Figure 6, spectrograms related to the LRU model appear similar to the target output than those of the S4D model, despite a discernible decrease in accuracy observed in Figure 7. In the case of the compressor, the capability to encode past information of the signal is critical, and S4D exhibits overall better accuracy. The ED and LRU models are designed to improve the temporal tracking abilities of basic LSTM, as confirmed by the results for both compressor devices. This improvement is particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters. These reach up to ms and seconds, respectively. In Figure 4 and 5, the difference across models is subtler, yet the LRU model consistently shows smaller RMS errors than other models and a more accurate match in the frequency domain as visible in Figure 6 and 7.
Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal. Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 6. However, the LRU model outperforms the others when considering the quantitative metrics. The LSTM model, in particular, tends to incorrectly predict excessive energy in the lower spectrum, as shown in Figure 7.
@@
| LA-2A | LSTM | CL 1B | LSTM | ||||
| ED | ED | ||||||
@@
| Helper | LSTM | Pultec | LSTM | ||||
| Saturator | ED | Passive | ED | ||||
| LRU | EQ | LRU | |||||
| S4D | S4D | ||||||
| Neutron’s | LSTM | ||||||
| Filter | ED | ||||||
| LRU | |||||||
| S4D | |||||||
@@
Generally, performance metrics are relatively close across models across most effects examined in this study. Although the S4D models slightly outperform other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study. The ED, LRU, and particularly the S4D models are more effective at encoding past signal information within their internal states. This is suggested by the case of the cases of optical compressors, where they significantly outperform other models. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters. Moreover, LSTM models tend to predict energy in the low-frequency spectrum more accurately.